Casinos use a six deck shoot for their tables. Find the probability of being dealt a blackjack an ace with one of those other face cards K,Q,J)
Answer #1

Number of cards in 6 decks = 6 * 52 = 312

Probabbility of getting two cards out of which one is Ace and other the face card = (24C1 * 72C1)/312C2

= (24*72)/(312*311/2)

= 3456/97032

= 0.0356 = 3.56% Answer